OpenStudy (brightemmerald):

Which of the binomials below is a factor of this trinomial? -2x^2 - 16x + 40 A. x - 5 B. x + 10 C. x - 10 D. x + 5

OpenStudy (mathstudent55):

Take a polynomial function f(x) of degree 2 or more. If a binomial x - a is a factor of f(x), then f(a) = 0.

OpenStudy (brightemmerald):

uuhhhh

OpenStudy (mathstudent55):

Now look at the options. For binomial x - 5, a = 5 For binomial x + 10, a = -10 Do the same for the other two options. Then evaluate the polynomial at each a. If the polynomial evaluates to zero, that is the factor.

OpenStudy (mathstudent55):

Here is another way of explaining this. Look at each option and compare it with x - a.

OpenStudy (brightemmerald):

i still do not understand

OpenStudy (mathstudent55):

A: x - 5 compare to x - a; a = 5 B: x + 10 compared to x - a: a = -10 C: x - 10 compared to x - a: x = 10 D: x + 5 compared to x - a: a = -5 Do you understand it so far?

OpenStudy (brightemmerald):

not really

OpenStudy (mathstudent55):

You need to have a binomial in the form x - a, where a number.

OpenStudy (mathstudent55):

You must have x followed by a subtraction, then a number.

OpenStudy (mathstudent55):

Look at option A: x - 5 Since you already have x followed by a subtraction, then the number 5, then for option A, which is x - 5, you can see that a = 5.

OpenStudy (mathstudent55):

Option B has x + 10. In this case, x is followed by an addition, but we need the form x - a, so we write x + 10 as x - (-10) This hows that in this case, a = -10

OpenStudy (mathstudent55):

We need each option to be in the form x - a A. x - 5 = x - a, so a = 5 B. x + 10 = x - (-10) = x - a, so a = -10 C. x - 10 = x - a, so a = 10 D. x + 5 = x - (-5) = x - a, so a = -5

OpenStudy (mathstudent55):

Now we evaluate the given polynomial at each value of a. If it evaluates to zero, then that is the option that is a factor.

OpenStudy (mathstudent55):

A. a = 5 -2x^2 - 16x + 40 = -2(5)^2 - 16(5) + 40 = -50 - 80 + 40 = -90 It does not evaluate to zero, so this is not it. B. a = -10 -2x^2 - 16x + 40 = -2(-10)^2 - 16(-10) + 40 = -200 + 160 + 40 = 0 It does evaluate to zero, so this is it. C. a = 10 -2x^2 - 16x + 40 = -2(10)^2 - 16(10) + 40 = -200 - 160 + 40 = -320 It does not evaluate to zero, so this is not it. D. a = -5 -2x^2 - 16x + 40 = -2(-5)^2 - 16(-5) + 40 = -50 + 80 + 40 = 70 It does not evaluate to zero, so this is not it.